OpenStudy (anonymous):

Simplify the expression. Show your work + (3^2 – 4^2)

TheSmartOne (thesmartone):

\(\sf\Large 3^2 = 3 \times 3 = ?\) \(\sf \Large4^2 = 4 \times 4 = ?\)